# Geometry-Aware Diffusion Models for Multiview Scene Inpainting

### v1

@@ 4.4 Ablation Studies
In Tab. 4, we ablate various inference-time strategies of our inpainting pipeline on scene completion. We observe that providing ground-truth depth maps has little effect on performance, highlighting the robustness of our method. However, ground-truth camera parameters have a more significant impact. This is mainly because optimizing camera parameters in DUSt3R involves complex global alignment, whereas, when camera parameters are known, optimizing the depth maps becomes a much simpler task.
In addition, for a wide-baseline dataset like NeRFiller, we find it is essential to run our autoregressive procedure, as a single reference cannot provide sufficient information for a wide baseline. The qualitative example in Fig. 10 also confirms our findings. Further ablation studies, including on geometric cues and training setups, are presented in §H.
SR 27.42 0.88 0.07 3.77 1232
AR 28.32 0.88 0.05 3.78 1235
AR 28.29 0.88 0.05 3.79 1231
AR 28.44 0.89 0.05 3.80 1252
AR 28.59 0.89 0.05 3.80 1250
Table 4: Ablation of available inputs and inpainting strategies on the NeRFiller scenes dataset. We denote the camera parameters as $\Pi$, depth maps as $D$, inpainting strategy as St., single-reference inpainting as SR, and autoregressive inpainting as AR. Note that the last row represents our base strategy.
@@ 5 Conclusion
In this paper, we proposed a novel approach to 3D scene inpainting task, without a 3D radiance field. While this avoids fusing in pixel space, which induces blurriness, it necessitates a novel way to fuse information across views via estimated scene geometry. We do so via training a diffusion-based inpainter, conditioned on appearance and geometric cues from a reference view-set, which enables fusion in the learned space of the generative model, thus retaining sharpness. The resulting multiview inpainting algorithm is highly versatile, capable of handling the sparse-view inpainting task, on which other methods struggle, and operating without cameras. We explored the efficacy of our method against two recent benchmarks, encompassing narrow and wide baseline 3D scenes, as well as the few-view scenario, showing state-of-the-art performance in all cases, in terms of both image quality and multiview consistency.
@@ A Overview
This document provides additional supplemental material to the main paper. §B details our proposed method, including scene geometry estimation, triangle mesh creation, the fusion of multiple reference images, training details, and inference details. §C outlines additional implementation details for our approach as well as the hyperparameters used. §D and §E describe details on the datasets and metrics used for evaluation, respectively. §F provides additional qualitative results. §G provides comprehensive results using the TSED metric. §H demonstrates additional ablations studies on various model design decisions. Finally, §I discusses the limitations of our work. Please see our project page for visualizations of results: https://geomvi.github.io.
@@ B.5 Parallel Process Fusion via Predicted Confidence Maps
3:  $\mathbf{C}^{\prime}_{b}=\mathbf{C}_{b}\,\land\,\neg\widehat{\mathbf{C}}_{f}$triangleright Backfaces but not front-faces. $\in\{0,1\}^{R\times H\times W}$
4:  $\widehat{\mathbf{C}}_{b}=\bigvee_{r}\mathbf{C}^{\prime}_{b}[r,:,:]$triangleright At least one backface exists. $\in\{0,1\}^{H\times W}$
5:  $\mathbf{C}^{\prime}_{s}=\mathbf{C}_{s}\,\land\,\neg(\widehat{\mathbf{C}}_{f}\lor\widehat{\mathbf{C}}_{b})$triangleright Shadows but no front-faces or backfaces. $\in\{0,1\}^{R\times H\times W}$
@@
Figure 7: Illustration of the fusion of multiple reference views using the confidence masks. We denote $I$ as the incomplete target image, $M$ as the inpainting mask, $\mathcal{R}$ as the set of reference images, $(A_{\mathcal{R}},\mathcal{G}_{\mathcal{R}})$ as reference-based appearance and geometric cues, $\Gamma$ as the fusion operator, $\mathbf{C}_{f},\mathbf{C}_{b},\mathbf{C}_{s}$ as the front-face, backface, and shadow confidence mask, $\mathcal{E}_{t}$ as the noise estimates at diffusion timestep $t$, $\widehat{\mathbf{F}}$ as the fused mask, and $\varepsilon_{t}$ as the fused noise estimate at diffusion timestep $t$, respectively. $\widehat{\mathbf{C}}_{f},\widehat{\mathbf{C}}_{b},\widehat{\mathbf{C}}_{s},\mathbf{C}^{\prime}_{b},\mathbf{C}^{\prime}_{s},\mathbf{C}_{\varnothing}$ represent intermediate variables of the fusion process; see Algorithm 1 for details. As shown, given a target image and a set of reference images, we first render the reference-based appearance and geometric cues, and then at each diffusion step, we fuse the noise estimates conditioned on different reference views using the predicted confidence masks.
@@ B.6 Training Details
Supervising Predicted Confidence Masks. We use the unperturbed meshes, $\mathcal{M}_{r}$, and $\mathcal{S}_{r}$, to render the ground-truth confidence masks. Specifically, we render $\mathcal{M}_{r}$ and $\mathcal{S}_{r}$ to obtain the unperturbed front-face, backface, and shadow masks, $\widehat{F}_{r},\widehat{B}_{r},\widehat{C}_{r}$. Given the sampled inpainting mask, $M$, the ground-truth front-face confidence mask is computed as
@@
where ,,$\land$,$\lor$,$\neg$ , ∨ , ¬ denote the logical “and”, “or”, and negation, respectively. This mask highlights the regions that are either outside the inpainting mask, or exclusively guided by front-facing surfaces (excluding the parts intersecting the shadow mask). The ground-truth backface confidence mask is computed as
@@ C Implementation Details
As mentioned in §3.2.3, we initialize our reference-guided inpainter as a Stable Diffusion v2, fine-tuned for inpainting [53, 60]. In training, we upweight Google Scanned Objects (GSO) so that the ratio of GSO to MS COCO samples in each epoch is 1/10. Since GSO lacks text captions, we use Kosmos-2 [45] to caption a front-facing view of each object. To make the model robust towards color and texture discrepancies across multiple views, we randomly augment the texture of the reference mesh using color jitter with a factor of 0.1 for brightness, contrast, saturation, and hue. To preserve classifier-free guidance [20] capabilities, we randomly drop the text prompt with a probability of 0.1. If the generated mask for an image is a 3D occluder volume, we randomly drop the reference content occluded by the occluder volume with a probability of 0.2. This will enable conditioning the model on not-yet-inpainted reference views. Notice that our model must be capable of both transferring reference information and also inpainting when no reference information is present. We use AdamW [35] with a learning rate of $10^{-4}$, weight decay of 0.010.010.01, $\beta_{1}=0.9$, and $\beta_{2}=0.999$. We train for approximately 8000 iterations, with a batch size of 96, and two gradient accumulation steps, on 16 NVIDIA L40 GPUs. The weight of the loss for the latent patches outside the inpainting mask is set to 0.10.10.1. During inference, we adopt the DDIM sampler [59] with 50 denoising steps.
@@ F Further Qualitative Results
In this section, we provide additional qualitative results. Fig. 8 visualizes examples of inpainted images’ depth maps from different inpainting methods. As shown, our method yields realistic geometry. Fig. 9 presents further qualitative results for the few-view inpainting task, confirming higher sharpness and visual plausibility than the baselines. Moreover, Fig. 10 shows a qualitative example of the ablation studies presented in §4.4, confirming ground-truth camera parameters and the autoregressive procedure affect the quality of the inpainted scene, while ground-truth depth maps have little impact. Finally, Fig. 11 shows a step-by-step example of autoregressive inpainting progress in the first autoregressive stage, inpainting a wide-baseline subset of the scene. As detailed in §3.3.3, we begin with a random view and select a wide-baseline subset of the scene, progressively inpainting the closest view from this subset to the already inpainted ones at each autoregressive step.
@@ G Comprehensive TSED Evaluation
Figure 10: A qualitative example from the ablation studies in §4.4 confirming ground-truth camera parameters and the autoregressive procedure affect the quality of the inpainted scene, whereas ground-truth depth maps have little impact.
@@
COCO + GSO Reprojected Hierarchical 29.25 0.89 0.05 3.72 1222
COCO + GSO Reprojected Closest camera 27.36 0.88 0.06 3.78 1204
COCO + GSO Reprojected Weighted average 29.45 0.89 0.06 3.77 1166
COCO + GSO Reprojected Single confidence 28.36 0.88 0.05 3.78 1223
@@
Table 5: Ablations of various design choices in training and fusion, including conditioning signals, datasets, and other algorithmic components. We denote the presence of geometric cues in the conditioning signals as $\mathcal{G_{\mathcal{R}}}$. ‘Closest camera’ and ‘Weighted average’ ignore the predicted confidence masks; the former solely conditions on the closest view that has been inpainted, and the latter takes a weighted average of the noise estimates, proportional to the inverse view distance between the reference view, $r$, and the target view, $t$, i.e., $\frac{1}{d((r,t))}$ (Eq. 1). ‘Single confidence’ means that, since backface and shadow masks are disabled, there is only one confidence signal, derived from the frontface mask. The first row represents our base model, which is superior on nearly all metrics, compared to other variants.
H Model Design Ablations
@@
In addition to ablating various inference-time strategies, presented in §4.4, we also ablate or vary several design decisions in our training and fusion strategies, including training datasets, availability of geometric cues ($\mathcal{G_{\mathcal{R}}}$) in the conditioning signals, whether to align the reference images to the coordinate of the target image by reprojection, and whether to use mesh perturbation. According to Tab. 5, we find that although mesh perturbation does not improve image-based metrics, it has a significant impact on video-based metrics, i.e., a higher image quality and consistency.
@@
Moreover, we find that it is essential to use our hierarchical fusion method, as alternative approaches such as “weighted average” and “closest camera” lead to lower performance. On the other hand, “weighted average” achieves the highest PSNR and SSIM among all settings. This is primarily because averaging multiple noise estimates may fuse inconsistent information, producing a blur artifact similar to NeRF renders. Since the inpainted images are already significantly blurred, fitting a NeRF does not introduce additional blurriness. This will result in higher consistency between the inpainted images and their corresponding NeRF renders, leading to higher PSNR and SSIM, though at the cost of lower overall image quality, as reflected in other metrics.
@@
We also observe the importance of conditioning the inpainter on the geometric cues. Note that in this case, fusion is performed using a single confidence mask; with backface and shadow masks disabled, the only confidence signal comes from the frontface mask.

### v2

@@ 4.4 Ablation Studies
In Tab. 4, we summarize our extensive ablation studies from §H, ablating key components of our inpainting pipeline on scene completion. We observe (first row) the importance of conditioning the inpainter on the geometric cues (§3.2.1). For wide-baseline datasets like NeRFiller, our autoregressive procedure (§3.3) is essential, since a single reference lacks sufficient information for a wide baseline (second row). Finally, we find that providing DUSt3R with ground-truth depth maps has minimal impact; however, known camera parameters significantly improve performance (third and fourth rows). This is mainly because optimizing the camera parameters in DUSt3R requires complex global alignment, while known cameras simplify depth optimization.
Variation PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ MUSIQ $\uparrow$ Corrs $\uparrow$
No Geometric Cues 28.36 0.88 0.05 3.78 1223
Single-Reference 27.42 0.88 0.07 3.77 1232
No GT Camera 28.29 0.88 0.05 3.79 1231
No GT Depth 28.44 0.89 0.05 3.80 1252
Full Model 28.59 0.89 0.05 3.80 1250
Table 4: Summary of key ablations for various design choices in the model and inpainting procedure on the NeRFiller dataset. Please see §4.4 for an explanation and §H for our exhaustive ablation studies.
@@ 5 Conclusion
In this paper, we introduced a novel approach to 3D scene inpainting task, without a 3D radiance field. While this avoids fusing in pixel space, which induces blurriness, it necessitates a novel way to fuse information across views via estimated scene geometry. We do so via training a diffusion-based inpainter, conditioned on appearance and geometric cues from a reference view-set, which enables fusion in the learned space of the generative model, thus retaining sharpness. The resulting multiview inpainting algorithm is highly versatile, capable of handling the sparse-view inpainting task, on which other methods struggle, and able to operate without camera poses. We explored the efficacy of our method on two recent benchmarks, encompassing narrow- and wide-baseline 3D scenes, as well as the few-view scenario, showing state-of-the-art performance in all cases, in terms of both image quality and multiview consistency.
@@ A Overview
This document provides additional supplemental material to the main paper. §B details our proposed method, including scene geometry estimation, triangle mesh creation, the fusion of multiple reference images, training details, and inference details. §C outlines additional implementation details for our approach as well as the hyperparameters used. §D and §E describe details on the datasets and metrics used for evaluation, respectively. §F provides additional qualitative results. §G provides comprehensive results using the TSED metric. §H demonstrates comprehensive ablation studies on various design decisions for our inpainting pipeline. Finally, §I discusses the limitations of our work. For more qualitative demonstrations, please see our project page: https://geomvi.github.io.
@@ B.5 Parallel Process Fusion via Predicted Confidence Maps
3:     $\mathbf{C}^{\prime}_{b}=\mathbf{C}_{b}\,\land\,\neg\widehat{\mathbf{C}}_{f}$triangleright Back-faces but not front-faces. $\in\{0,1\}^{R\times H\times W}$
4:     $\widehat{\mathbf{C}}_{b}=\bigvee_{r}\mathbf{C}^{\prime}_{b}[r,:,:]$triangleright At least one back-face exists. $\in\{0,1\}^{H\times W}$
5:     $\mathbf{C}^{\prime}_{s}=\mathbf{C}_{s}\,\land\,\neg(\widehat{\mathbf{C}}_{f}\lor\widehat{\mathbf{C}}_{b})$triangleright Shadows but no front-faces or back-faces. $\in\{0,1\}^{R\times H\times W}$
@@
Figure 7: Illustration of the fusion of multiple reference views using the confidence masks. We denote $I$ as the incomplete target image, $M$ as the inpainting mask, $\mathcal{R}$ as the set of reference images, $(A_{\mathcal{R}},\mathcal{G}_{\mathcal{R}})$ as reference-based appearance and geometric cues, $\Gamma$ as the fusion operator, $\mathbf{C}_{f},\mathbf{C}_{b},\mathbf{C}_{s}$ as the front-face, back-face, and shadow confidence mask, $\mathcal{E}_{t}$ as the noise estimates at diffusion timestep $t$, $\widehat{\mathbf{F}}$ as the fused mask, and $\varepsilon_{t}$ as the fused noise estimate at diffusion timestep $t$, respectively. $\widehat{\mathbf{C}}_{f},\widehat{\mathbf{C}}_{b},\widehat{\mathbf{C}}_{s},\mathbf{C}^{\prime}_{b},\mathbf{C}^{\prime}_{s},\mathbf{C}_{\varnothing}$ represent intermediate variables of the fusion process; see Algorithm 1 for details. As shown, given a target image and a set of reference images, we first render the reference-based appearance and geometric cues, and then at each diffusion step, we fuse the noise estimates conditioned on different reference views using the predicted confidence masks.
@@ B.6 Training Details
Supervising Predicted Confidence Masks. We use the unperturbed meshes, $\mathcal{M}_{r}$, and $\mathcal{S}_{r}$, to render the ground-truth confidence masks. Specifically, we render $\mathcal{M}_{r}$ and $\mathcal{S}_{r}$ to obtain the unperturbed front-face, back-face, and shadow masks, $\widehat{F}_{r},\widehat{B}_{r},\widehat{C}_{r}$. Given the sampled inpainting mask, $M$, the ground-truth front-face confidence mask is computed as
@@
where ,,$\land$,$\lor$,$\neg$ , ∨ , ¬ denote the logical “and”, “or”, and negation, respectively. This mask highlights the regions that are either outside the inpainting mask, or exclusively guided by front-facing surfaces (excluding the parts intersecting the shadow mask). The ground-truth back-face confidence mask is computed as
@@ C Implementation Details
As mentioned in §3.2.3, we initialize our reference-guided inpainter as a Stable Diffusion v2, fine-tuned for inpainting [55, 63]. In training, we upweight Google Scanned Objects (GSO) so that the ratio of GSO to MS COCO samples in each epoch is 1/10. Since GSO lacks text captions, we use Kosmos-2 [46] to caption a front-facing view of each object. To make the model robust towards color and texture discrepancies across multiple views, we randomly augment the texture of the reference mesh using color jitter with a factor of 0.1 for brightness, contrast, saturation, and hue. To preserve classifier-free guidance [20] capabilities, we randomly drop the text prompt with a probability of 0.1. If the generated mask for an image is a 3D occluder volume, we randomly drop the reference content occluded by the occluder volume with a probability of 0.2. This will enable conditioning the model on not-yet-inpainted reference views. Notice that our model must be capable of both transferring reference information and also inpainting when no reference information is present. We use AdamW [36] with a learning rate of $10^{-4}$, weight decay of $0.01$, $\beta_{1}=0.9$, and $\beta_{2}=0.999$. We train for approximately 8000 iterations, with a batch size of 96, and two gradient accumulation steps, on 16 NVIDIA L40 GPUs. The weight of the loss for the latent patches outside the inpainting mask is set to $0.1$. During inference, we adopt the DDIM sampler [62] with 50 denoising steps.
@@ F Further Qualitative Results
In this section, we provide additional qualitative results.
F.1 Comparison of Depth Maps
Fig. 8 visualizes examples of inpainted images’ depth maps from different inpainting methods. As shown, our method yields realistic geometry.
F.2 Few-View Inpainting
Fig. 9 presents further qualitative results for the few-view inpainting task, confirming higher sharpness and visual plausibility than the baselines.
F.3 An Example of Autoregressive Inpainting
Fig. 10 shows a step-by-step example of autoregressive inpainting progress in the first autoregressive stage, inpainting a wide-baseline subset of the scene. As detailed in §3.3.3, we begin with a random view and select a wide-baseline subset of the scene, progressively inpainting the closest view from this subset to the already inpainted ones at each autoregressive step.
F.4 Qualitative Comparison with MALD-NeRF
Fig. 11 presents a direct qualitative comparison to MALD-NeRF [31], the current state of the art. Here, we highlight the inconsistencies in MALD-NeRF’s outputs, primarily caused by common NeRF artifacts, such as floaters or flawed geometry, which compromise geometric plausibility. As a result, the artifacts in MALD-NeRF prevent SIFT from detecting sufficient high-quality feature correspondences, leading to a lower TSED metric.
@@ G Comprehensive TSED Evaluation
Figure 11: Qualitative comparison of our method with MALD-NeRF. Each inpainted view is accompanied by a zoomed-in version on the sides, highlighting inconsistencies.
@@
H Comprehensive Ablation Studies
We extend the ablation studies presented in §4.4 along three key axes: (i) comparison with the naive baseline of independent 2D inpainting, (ii) ablation of inference-time strategies, and (iii) ablating or varying various design choices in our training and fusion strategies.
H.1 Comparison with Independent 2D inpainting
Method PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ MUSIQ $\uparrow$ Corrs $\uparrow$
Stable Diffusion (2D) [55] 24.69 0.85 0.10 3.77 1120
ControlNet (2D) [85] 21.33 0.83 0.14 3.70 1024
BrushNet (2D) [24] 22.84 0.83 0.13 3.77 1081
Ours 28.59 0.89 0.05 3.80 1250
Table 5: Evaluating our method against independent inpainting on scene completion. “2D” indicates independent 2D inpainting.
As observed in prior work [75], independent inpainting fails to produce consistent content across views. Since reference-based geometry-awareness is one of the core components of our approach, we also present a comparison between our geometry-aware inpainting and the naive baseline of geometry-unaware inpainting (i.e., independent 2D inpainting), in Tab. 5. We compare our approach to three state-of-the-art diffusion-based 2D inpainting methods: Stable Diffusion [55] (which our model is based on), ControlNet [85], and BrushNet [24]. Note that our model, just as for the 2D inpainter baselines, is also a latent diffusion model with a similar architecture, operating on a single image at a time. In other words, we do not use an explicit or implicit 3D radiance field when inpainting the views. The difference lies only in the conditioning signals: our diffusion model is informed by the 3D world and other views through the various cues passed to the generator at inference time. After inpainting all the views, similar to our method, a NeRF is fit to the inpainted views and the rendered images and videos are assessed to compute the evaluation metrics (refer to §4.2 for details). As demonstrated, our method significantly outperforms all baselines, confirming its superior quality in the context of 3D inpainting.
H.2 Ablation of Inference-Time Strategies
SR 27.42 0.88 0.07 3.77 1232
AR 28.32 0.88 0.05 3.78 1235
AR 28.29 0.88 0.05 3.79 1231
AR 28.44 0.89 0.05 3.80 1252
AR 28.59 0.89 0.05 3.80 1250
Table 6: Ablation of available inputs and inpainting strategies on the NeRFiller scenes dataset. We denote the camera parameters as $\Pi$, depth maps as $D$, inpainting strategy as St., single-reference inpainting as SR, and autoregressive inpainting as AR. Note that the last row represents our full strategy.
Figure 15: A qualitative example from the ablation of inference-time strategies, confirming ground-truth camera parameters and the autoregressive procedure affect the quality of the inpainted scene, whereas ground-truth depth maps have little impact.
@@ H.2 Ablation of Inference-Time Strategies
COCO + GSO Reprojected Single confidence 28.36 0.88 0.05 3.78 1223
COCO + GSO Reprojected Weighted average 29.45 0.89 0.06 3.77 1166
COCO + GSO Reprojected Closest camera 27.36 0.88 0.06 3.78 1204
COCO + GSO Reprojected Hierarchical 29.25 0.89 0.05 3.72 1222
@@
Table 7: Ablation of various design choices in training and fusion, including conditioning signals, datasets, and other algorithmic components. We denote the presence of geometric cues in the conditioning signals as $\mathcal{G_{\mathcal{R}}}$. ‘Closest camera’ and ‘Weighted average’ ignore the predicted confidence masks; the former solely conditions on the closest view that has been inpainted, and the latter takes a weighted average of the noise estimates, proportional to the inverse view distance between the reference view, $r$, and the target view, $t$, i.e., $\frac{1}{d((r,t))}$ (Eq. 1). ‘Single confidence’ means that, since back-face and shadow masks are disabled, there is only one confidence signal, derived from the front-face mask. The last row represents our full model, which is superior on nearly all metrics, compared to other variants.
In Tab. 6, we ablate various inference-time strategies of our inpainting pipeline on scene completion. We observe that, for wide-baseline datasets like NeRFiller, our autoregressive procedure (§3.3) is essential, as a single reference lacks sufficient information for a wide baseline (first row). We also find that providing DUSt3R with ground-truth depth maps has little effect on performance, highlighting the robustness of our method. However, ground-truth camera parameters have a more significant impact (second to fourth rows). This is mainly because optimizing camera parameters in DUSt3R involves complex global alignment, whereas, when camera parameters are known, optimizing the depth maps becomes a much simpler task. The qualitative example in Fig. 15 also confirms our findings.
H.3 Model Design Ablation
@@
Finally, we ablate or vary several design decisions in our training and fusion strategies, including training datasets, availability of geometric cues ($\mathcal{G_{\mathcal{R}}}$) in the conditioning signals, whether to align the reference images to the coordinate of the target image by reprojection, and whether to use mesh perturbation. According to Tab. 7, we find that although mesh perturbation (§3.2.3) does not improve image-based metrics, it has a significant impact on video-based metrics, i.e., a higher image quality and consistency.
@@
Moreover, we find that it is essential to use our hierarchical fusion method (§3.3.1), as alternative approaches such as “weighted average” and “closest camera” lead to lower performance. On the other hand, “weighted average” achieves the highest PSNR and SSIM among all settings. This is primarily because averaging multiple noise estimates may fuse inconsistent information, producing a blur artifact similar to NeRF renders. Since the inpainted images are already significantly blurred, fitting a NeRF does not introduce additional blurriness. This will result in higher consistency between the inpainted images and their corresponding NeRF renders, leading to higher PSNR and SSIM, though at the cost of lower overall image quality, as reflected in other metrics.
@@
We also observe the importance of conditioning the inpainter on the geometric cues (§3.2.1). Note that in this case, fusion is performed using a single confidence mask; with back-face and shadow masks disabled, the only confidence signal comes from the front-face mask.